# Send Message to the Future? Blockchain-based Time Machines for Decentralized Reveal of Locked Information

### v2

@@ Abstract
Conditional information reveal systems automate the release of information upon meeting specific predefined conditions, such as time or location. This paper introduces a breakthrough in the understanding, design and application of conditional information reveal systems that are highly secure and decentralized. By designing a new practical timed-release cryptography system and a verifiable secret sharing scheme, a novel data sharing system is devised on the blockchain that ‘sends messages in the future’ with highly accurate decryption times. This paper provides a complete evaluation portfolio of this pioneering paradigm, including analytical results, a validation of its robustness in the Tamarin Prover and a performance evaluation of a real-world, open-source system prototype deployed across the globe. Using real-world election data, we also demonstrate the applicability of this innovative system in e-voting, illustrating its capacity to secure and ensure fair electronic voting processes.
@@ I Introduction
The advent of the digital age has made information sharing and access highly complex and sensitive. Amid these complexities, conditional information reveal stands as a compelling problem area of exploration. Conditional information reveal uses computer systems to automatically reveal information upon the verification of certain requirements. From releasing information after a certain point of time [1, 2], to disseminating information based on geographical location [3, 4, 5] and restricting access to classified documents based on identity [6, 7], these are all instances of information reveal dictated by a certain condition of time, location, or access right. One such condition, pivotal in its role across a multitude of applications, is time.
The study of time-based conditional information reveal systems is known as timed-release cryptography, or time lock encryption. As a cryptographic technique, it guarantees the confidentiality of a message until a specific point in time, upon which the message is automatically decrypted and becomes accessible to the intended parties. Despite its practical potential, its journey from theory to widespread usage has been hindered by the need to ensure accurate decryption time, message confidentiality against adversarial parties, and energy efficiency [1, 2].
This paper overcomes these challenges by modelling, for the first time, time lock encryption as an automated conditional information reveal construction, and introduces a novel architecture designed to achieve conditional information reveal in a decentralized way. Building upon this new architecture, this paper further proposes a novel timed-release cryptography scheme using time as the condition for information reveal.
Backed by blockchain technology and threshold cryptography, the proposed system allows a consortium of distributed secret holders to safeguard sensitive information of clients with a minimum level of trust. The decentralization, immutability, and transparency nature of blockchains ensures authenticity, availability and assurance in the communication among clients and secret holders. Furthermore, a blockchain-based incentive mechanism is applicable to ensure secret holders’ honesty, thereby enhancing the overall system reliability.
Notably, this paper also proposes a fully verifiable secret sharing scheme when constructing the timed-release cryptography system to enhance its efficiency and security. Unlike the common definition of verifiable secret sharing (VSS) in existing literature [8] that only provides verifiability of the generation and dissemination of secret shares, This novel secret sharing also provides verifiability of the reconstruction of the secret with low communication costs.
@@
  • A blockchain-based architecture to achieve conditional information reveal in a decentralized way, which provides a high level of security and reliability to conditional information reveal service providers such as election authorities. It uses smart contracts to automate fair rewards distribution to multiple secret holders.
  • A novel fully verifiable secret sharing scheme to achieve timed-release cryptography under the proposed architecture to provide verifiability in secret reconstruction and ensure efficient communication under the blockchain communication model.
@@
The rest of the paper is structured as follows: Section 2 reviews the related work in conditional information reveal systems and timed-release cryptography. Section 3 introduces conditional information reveal systems with a novel paradigm and architecture. Section 4 illustrates how the architecture can be used to construct a novel timed-release cryptography system. Section 5 presents a new cryptographic protocol that secures the proposed system. Section 6 shows the advantages of the proposed system compared to the existing methods. Section 7 showcases the applicability of the proposed system in an e-voting scenario and Section 8 concludes the paper with an outline of future work.
@@ II-A1 The Time Lock Puzzles Approach
In the time lock puzzles approach, the answer to the computational puzzle serves as the decryption key to the message [12, 13]. The computational difficulty of the puzzle is set to adjust the expected puzzle solving time given an estimation of available computational resources.
@@
In recent years, the emergence of blockchain technology has advanced research in constructing time lock puzzles. The proof-of-work consensus mechanism [21] offers a new way to construct time lock puzzles with more stable puzzle-solving times. Inspired by proof-of-work blockchains, researchers have studied the challenge of dynamically controlling puzzle difficulty by finding a “computational reference clock” [12]. This clock measures the mapping between computational difficulty and real-world time. For example, Bitcoin is designed to map a puzzle to ten minutes real-world time by adjusting its difficulty according to the available computational resources. Various methods are proposed in existing work to convert the proof-of-work consensus to a time lock puzzle system [22, 12, 23].
While these proposals are theoretically feasible, they still face practical limitations. First, the puzzle solving time is not stable enough to secure real-world time-sensitive messages as computational puzzles always involve random searches.
@@
As the largest computational reference clock in the world that aims to solve a puzzle every ten minutes, the reality is far from its target. According to the Bitcoin block data from the July 2022 to July 2023 provided by Google BigQuery [24], although the average block time is computed as around 9.8 minutes, which is close to the target of 10 minutes per block, Figure 1 indicates that creation time for individual blocks is not stable. The chance of a new block being created in 1 minute is more than two times larger than being created in exactly 10 minutes.
@@ II-A2 The Secret Holders Approach
Centralized designs [1, 25, 26, 27, 28, 29, 30] rely on one independent party acting as a secret holder. These methods focus on cryptographic level security under the assumption of the honesty and activeness of the single secret holder. Security and robustness at a system level are not considered, such as potential single point of failure, high level of trust required for the secret holder. The decentralized methods [31, 2, 32, 33, 34] utilize threshold cryptography to provide fault tolerance while involving multiple secret holders to guard a secret.
Although trust is distributed to multiple holders in decentralized solutions, there are yet security problems in these systems. The first problem is the lack of reliability due to a lack of incentives. The question of “why should clients trust the honesty of secret holders” is not addressed in the existing proposals. Nothing stops secret holders from cheating. The second problem is data availability and authenticity. The need for “public bulletin board” [32] or “publicly available location” to store data [2] is documented in existing proposals, but the robustness and authentication are not explicitly addressed. The third problem is the lack of counter-measurements to adversarial secret holders in the system. Existing proposals do not consider the probability of having adversarial holders in the system, lacking measurements to prevent, detect and dispute adversarial holders.
@@
The idea of using secret holders to guard information is expected to introduce security challenges related to honesty and robustness of the secret holders. Therefore, we provide a comparative analysis of the proposed method against existing methods in Table I, focusing on several critical security features. We evaluate each approach based on its resistance to single points of failure, incentives for honesty among participants, resilience to time source poisoning (changing secret holders local time by attacking the time servers), resilience to Sybil attacks, and verifiability of released secrets.
TABLE I: Security Comparison of Different Approaches
Method Single point of failure resistance Incentives for honesty Time source poisoning resistance Sybil attack resistance Verifiability of released secret
@@
Decentralized on-demand: [2] Yes No No No No
@@
The decentralized on-demand approach only provides services upon receiving clients’ requests. The existing proposal [2] is closest to our approach as our approach also falls into this category. However, it does not offer incentives for honesty, resistance to time source poisoning, or Sybil attack resistance. It also falls short in terms of verifiability, highlighting the need for an approach that addresses these shortcomings.
Presented in the remaining sections, our approach distinguishes itself by leveraging security features from the blockchain and incorporating a verifiable cryptographic protocol, providing security features that are not achieved in other existing approaches.
@@ II-B Secret Sharing Schemes for Secret Holders Approach
The original Shamir’s Secret Sharing scheme does not provide verifiability to the correctness of secret shares [35]; existing work on verifiable secret sharing [36, 37] only provides verifiability in the distribution phase to ensure that holders receive correct shares. However, to the best of our knowledge, verifiability that ensures holders submit correct shares has not been demonstrated. In the case of timed-release cryptography, this introduces attack opportunities for adversarial holders to submit incorrect shares or for adversarial clients to make malicious claims that holders submit incorrect shares.
Moreover, secret sharing schemes often require peer-to-peer communication channels to distribute secret shares. When applied to timed-release cryptography, clients send secret shares to secret holders and also the time to reveal the shares. It is hard to ensure each secret holder receives the correct secret share and the same reveal time via a peer-to-peer channel. However, applying existing secret sharing schemes in a public communication channel results in significant communication overhead, given that clients are required to publish all encrypted shares to deliver a unique secret share to each secret holder. This results in a large message size of $n\cdot|s|$, where $n$ is the number of holders and $|s|$ is the size of a secret share. Each holder receives all $n$ encrypted secret shares, even though only one share is relevant to each holder.
Consequently, a new cryptographic protocol is introduced to achieve verifiability in the recovery phase and reduce communication costs.
@@ II-B1 Required Background
III Conditional Information Reveal Systems
This section introduces a generic automated conditional information reveal system with the aim to propose a paradigm for such systems and an architecture for decentralized conditional information reveal based on blockchain and smart contracts. The paradigm and architecture provides a strong foundation to the modeling of timed-release cryptography systems.
III-A Requirements for a New Paradigm
While conditional information reveal systems exhibit certain similarities, their design is confronted by distinct challenges. For instance, as a widely used identity-based conditional information reveal system, a foundational presumption in prevailing single-sign-on (SSO) systems is the implicit trust required between service providers and clients towards authentication service providers. Consequently, its centralized architecture introduces vulnerabilities, most notably, a single point of failure. In the realm of location verification, it is challenging to identify a method that is simultaneously reliable, efficient, and precise for positioning. Despite the prevalence of conditional information reveal systems, a systematic approach to identifying the challenges they face is absent. Therefore, this paper introduces a new paradigm to systematically study a broad spectrum of such systems.
In summarizing the core components of conditional information reveal systems, we identify two roles, three stages, and four challenges that characterize these systems. The two roles are clients who request a specific service and information holders who handle requests and reveal information upon satisfaction of conditions, for instance, time, location or identity. The three stages for clients and holders are request, condition check, and information reveal. Clients initiate the process by submitting requests to information holders, specifying the conditions for information reveal. The holders monitor the conditions and reveal the information upon satisfaction of the requirements. In some systems, as an incentive mechanism, information holders may receive rewards for successfully executing tasks as per client requests.
Constructing a conditional information reveal system poses challenges at each step of the process. These key challenges include:
@@ III-A Requirements for a New Paradigm
This observation of challenges helps to identify problems with existing timed-release cryptography proposals. While the existing proposals primarily concentrate on refining cryptographic protocols, these challenges, at the systemic level, remain unsolved. For instance, the distributed secret holder approach necessitates both a public bulletin board and a peer-to-peer communication channel. However, discussions on their implementation and reliability are missing. Similarly, using time as a condition for information reveal might appear straightforward, yet the critical issue of maintaining synchronized time across multiple holders has not received sufficient attention.
Moreover, a trade-off is evident in existing proposals regarding the verifiability of revealed information and providing incentives to keep secret holders active. Some proposals [34, 32, 33] achieve verifiable request completion by regularly supplying timed-release key pairs, irrespective of client demand. This approach, while ensuring verifiability, does not incentivize secret holders to remain active, as it requires holders to operate continuously, even in the absence of client requests. Conversely, the on-demand secret reveal approach [2], though more responsive to actual demand, lacks mechanisms to verify the correctness of the information once revealed.This trade-off highlights the need for a new method to provide strong security by ensuring both incentivization and verifiability.
III-B A Decentralized Architecture for Conditional Information Reveal
Ensuring the required conditions are securely met to share the information is challenging, especially when the systems are security and privacy sensitive. In such scenarios, it becomes difficult to rely on a single information holder to complete the task, as it requires a high level of trust to a single party and is vulnerable to single point of failure. Involving multiple holders can improve system reliability.
By leveraging the security features of blockchain systems, smart contracts facilitate reliable communication between clients and holders within a conditional information reveal system. Additionally, smart contracts handle escrow functionality, allowing clients to deposit rewards for holders when initiating a request, which can be automatically distributed by the contract code upon completion. The use of smart contracts as a communication platform ensures secure and transparent interactions, forming the backbone of the system.
@@ III-B A Decentralized Architecture for Conditional Information Reveal
Figure 2 illustrates the modules in the proposed decentralized conditional information reveal architecture. This generic architecture provides adaptability to different conditional information reveal systems, accommodating various information sharing conditions. Computation and verification of secret shares are functions executed locally by holders; holders communicate and self-organize the network through a smart contract.
III-B1 Blockchain as a Communication Channel
The adoption of blockchain technology in the conditional information reveal framework introduces a new way of communication compared to traditional off-chain environments. In this context, information holders and clients interact with smart contracts deployed on a blockchain network. This form of communication offers key advantages such as immutability and non-repudiation of messages, which are crucial for maintaining accountability and ensuring honest behavior among all parties involved in the system.
However, this blockchain-based communication model also presents challenges in scenarios where privacy-sensitive information is involved [40]. Since all messages within the blockchain are publicly accessible to all participants, including requests from clients and information shared by secret holders, the confidentiality of sensitive information may be compromised. This limitation needs to be addressed, particularly for privacy-sensitive conditional information reveal systems.
In subsequent sections, this paper proposes a new cryptographic protocol that addresses the challenge of efficiently sharing secrets through blockchain as a public communication channel. This protocol aims to accommodate the communication model, while preserving the privacy and security of sensitive information in time-based decentralized conditional information reveal systems.
III-B2 Network Autonomy
In a decentralized conditional information reveal system, multiple secret holders remain honest and active to serve requests and play a crucial role in governing the network and protecting it from malicious holders. This decentralization eliminates the need for external parties to manage the network, making it self-governing. Network autonomy is achieved by disputing secret holders and using protocols for verifiable information sharing.
To ensure network security, secret holders can dispute the actions of other holders via the smart contract. Sufficient disputes against a secret holder results in blacklisting and lost of cryptocurrencies as a punishment. This self-policing mechanism relies on the assumption that a majority of secret holders are honest and actively monitor the network.
Verifiability of information shared by holders is essential for raising disputes. The specific protocols for achieving verifiable information sharing may vary depending on the scenario. Designing a decentralized conditional information reveal system under this architecture includes protocols with which holders verify the information shared by their peers. In case of disputes or claims by malicious clients, this verification process serves as proof that a task is not correctly completed.
With network autonomy, the network becomes self-governing and can effectively mitigate risks from malicious holders and false claims from clients.
III-B3 Incentives and Cryptoeconomics
In a trustless setting, the assumption that a majority of secret holders are honest in a decentralized conditional information reveal system is realistic by providing incentives to secret holders. With blockchain as the underlying platform, the system incorporates cryptocurrencies seamlessly. Therefore, a cryptoeconomic system can be made available for conditional information reveal to ensure secret holders’ honesty with flows of rewards.
Clients incentivize holders by rewarding them with cryptocurrencies for executing their requests. This is a financial incentive for secret holders to perform their tasks diligently and efficiently. By receiving rewards for their services, secret holders are motivated to act honestly and fulfil their obligations to the clients. The payment for the service employs a specific token as the currency (such as the ERC-20 token in Ethereum [41]). These incentives are designed to strike a balance between secret holders’ profitability and clients’ service requests. Clients acquire these tokens as payment for accessing the service. The received tokens by secret holders can later be sold for profits. As service demand surges, the token value escalates, and conversely, it diminishes when demand dips.
In addition to client payments, secret holders are also required to deposit cryptocurrencies when joining the network. This deposit acts as a security measure and can be used as a form of collateral. In the event that a secret holder is identified as malicious or behaves dishonestly, their deposit can be used as a reward for other holders, who report the malicious behavior. This creates an additional layer of incentives for secret holders to actively monitor the network, report malicious activities, and maintain the overall security and integrity of the system.
The use of smart contracts simplifies the implementation of these incentive mechanisms. Smart contracts provide the necessary infrastructure to handle the deposit, escrow the funds, and automatically distribute the rewards to secret holders based on predefined rules. By using cryptocurrencies and smart contracts in a decentralized conditional information reveal system, the incentive system becomes efficient, transparent, and highly automated.
IV Time-based Conditional Information Reveal
This section illustrates a timed-release cryptography system using the proposed decentralized conditional information reveal architecture. The proposed system addresses the key challenges of conditional information reveal system design, resulting in a system that achieves higher security and reliability.
IV-A Applying the Blockchain-based Architecture

### v3

@@ Abstract
Conditional information reveal systems automate the release of information upon meeting specific predefined conditions, such as time or location. This paper introduces a breakthrough in the understanding, design, and application of conditional information reveal systems that are highly secure and decentralized. By designing a new practical timed-release cryptography system and a secret sharing scheme with reveal-verifiability, a novel data sharing system is devised on the blockchain that ‘sends messages in the future’ with highly accurate decryption times. Notably, the proposed secret sharing scheme applies to other applications requiring verifiability of revealed secret shares. This paper provides a complete evaluation portfolio of this pioneering paradigm, including analytical results, a validation of its robustness in the Tamarin Prover and a performance evaluation of a real-world, open-source system prototype deployed across the globe. Using real-world election data, we also demonstrate the applicability of this innovative system in e-voting, illustrating its capacity to secure and ensure fair electronic voting processes.
@@ I Introduction
Conditional information reveal uses computer systems to automatically reveal information upon the verification of certain requirements. From releasing information after a certain point of time [1, 2], to disseminating information based on geographical location [3, 4, 5] and restricting access to classified documents based on identity [6, 7], these are all instances of information reveal dictated by a certain condition of time, location, or access right. One such condition, pivotal in its role across a multitude of applications, is time.
The study of time-based conditional information reveal systems is known as timed-release cryptography or time lock encryption. It guarantees the confidentiality of a message until a specific point in time, upon which the message is automatically decrypted and becomes accessible to the intended parties. Despite its practical potential, its journey from theory to widespread usage has been hindered by the need to ensure accurate decryption time, message confidentiality against adversarial parties, and energy efficiency [1, 2].
This paper introduces a smart-contract-based timed-release cryptography system with improvements in security and efficiency compared to the existing proposals [8, 9, 10, 11, 12]. Moreover, this paper shows how general conditional information reveal systems can be secured using the same architecture as the proposed timed-release cryptography system.
Smart contract and threshold cryptography are used in the proposed system to allow a consortium of distributed secret holders to safeguard sensitive information with a minimum level of trust. The decentralization, immutability, and transparency nature of blockchains ensures authenticity, availability and assurance in the communication among clients and secret holders. Furthermore, a blockchain-based incentive mechanism is applicable to ensure secret holders’ honesty, thereby enhancing the overall system reliability.
Notably, this paper also proposes a secret sharing scheme with reveal-verifiability for conditional information reveal systems to enhance their efficiency and security. The common definition of verifiable secret sharing (VSS) and publicly verifiable secret sharing (PVSS) in existing literature [13] only provides verifiability of the generation and dissemination of secret shares. However, in timed-release cryptography, verifiability of the revealed secret shares is also required to ensure the reconstructed secret is the one sent by the client. This novel secret sharing scheme introduces this property named reveal-verifiability, that provides public verifiability in the secret reconstruction stage.
@@
  • A smart-contract-based architecture to achieve secure conditional information reveal in a decentralized way. A secure timed-release cryptography system is constructed under this architecture.
  • A novel secret sharing scheme with reveal-verifiability to provide verifiability in secret reconstruction and efficient communication under a public communication model.
@@
The rest of the paper is structured as follows: Section 2 reviews the related work in conditional information reveal systems and timed-release cryptography. Section 3 introduces a secure architecture for conditional information reveal systems using smart contracts and secret sharing. Section 4 presents a smart-contract-based timed-release cryptography system. Section 5 presents a new cryptographic protocol that secures the proposed system. Section 6 shows the advantages of the proposed system compared to the existing methods. Section 7 showcases the applicability of the proposed system in an e-voting scenario and Section 8 concludes the paper with an outline of future work.
@@ II-A1 The Time Lock Puzzles Approach
In the time lock puzzles approach, the answer to the computational puzzle serves as the decryption key to the message [17, 18]. The computational difficulty of the puzzle is set to adjust the expected puzzle-solving time given an estimation of available computational resources.
@@
In recent years, the emergence of blockchain technology has advanced research in constructing time lock puzzles. The proof-of-work consensus mechanism [26] offers a new way to construct time lock puzzles with more stable puzzle-solving times. Inspired by proof-of-work blockchains, researchers have studied the challenge of dynamically controlling puzzle difficulty by finding a “computational reference clock” [17]. This clock measures the mapping between computational difficulty and real-world time. For example, Bitcoin is designed to map a puzzle to ten minutes of real-world time by adjusting its difficulty according to the available computational resources. Various methods are proposed in existing work to convert the proof-of-work consensus to a time lock puzzle system [27, 17, 28, 29, 30, 31].
Existing works try to produce a more accurate mapping between puzzle difficulty and the required time to solve the puzzle, for example, by using verifiable delay functions (VDF) to create “proof of sequential work” systems [29, 30, 31] that reduces the randomness involved in the puzzle solving time. However, to the best of our knowledge, there are no studies to improve the puzzle difficulty adjustment mechanism, which is found to be problematic in the proof-of-work consensus when requiring a stable puzzle-solving time.
@@
As the largest computational reference clock in the world that aims to solve a puzzle every ten minutes, the reality is far from its target. According to the Bitcoin block data from July 2022 to July 2023 provided by Google BigQuery [32], although the average block time is computed as around 9.8 minutes, which is close to the target of 10 minutes per block, Figure 1 indicates that creation time for individual blocks is not stable. The chance of a new block being created in 1 minute is more than two times larger than being created in exactly 10 minutes. Although with recent works on proof of sequential work, a better distribution can be achieved, it is still difficult to have a robust and dynamic difficulty adjustment mechanism to adjust the puzzle difficulties with the available computing powers. Current proof-of-work consensus adjusts puzzle difficulties based on the average block time of the previous 2016 blocks. When applied to time lock encryption, a malicious party can decide to use less or more computing power to slow down or speed up solving a particular puzzle to affect specific time lock messages with little effect on the difficulties of subsequent puzzles.
@@ II-A2 The Secret Holders Approach
Centralized designs [1, 33, 34, 35, 36, 37, 38] rely on one independent party acting as a secret holder. These methods focus on cryptographic-level security under the assumption of the honesty and activeness of the single secret holder. Security and robustness at a system level are not considered, such as potential single point of failure, and high level of trust required for the secret holder. The decentralized methods [39, 2, 40, 41, 42] utilize threshold cryptography to provide fault tolerance while involving multiple secret holders to guard a secret.
Although trust is distributed to multiple holders in decentralized solutions, there are yet security problems in early secret holder proposals. The first problem is the lack of reliability due to a lack of incentives. The question of “why should clients trust the honesty of secret holders” is not addressed in the early proposals. Nothing stops secret holders from cheating. The second problem is data availability and authenticity. The need for “public bulletin board” [40] or “publicly available location” to store data [2] is documented in existing proposals, but the robustness and authentication are not explicitly addressed.
The development of blockchains and smart contracts addresses the challenges of incentives, data availability and authenticity. Recent proposals [8, 9, 10, 11, 12] use smart contracts that integrate cryptocurrencies to provide incentives for secret holders, and ensure secret shares are immutable and available after revealing. However, none of these proposals provides undeniable reveal-verifiability, that is, secret shares revealed by honest secret holders can be claimed as incorrect shares in the presence of malicious clients. The possibility of having malicious clients that frame secret holders is only mentioned by Ning et al. [9], but a solution to prevent false claims of incorrect shares is not proposed. Other proposals [8, 10, 11, 12] do not consider malicious clients in their security models.
@@
We provide a comparative analysis of the proposed method against existing secret holder methods in Table I, focusing on several critical security features. Each approach is evaluated based on the following perspectives: resistance to single points of failure, incentives for honesty among participants, resilience to time source poisoning (changing secret holders local time by attacking the time servers), resilience to Sybil attacks, and reveal-verifiability under malicious clients.
TABLE I: Security Comparison of Different Secret Holder Approaches
Method Single point of failure resistance Incentives for honesty Time source poisoning resistance Sybil attack resistance Reveal-verifiability under malicious clients
@@
Decentralized on-demand (off-chain): [2] Yes No No No No
Decentralized on-demand (smart-contract-based): [8, 9, 10, 11, 12] Yes Yes No Yes No
@@
The decentralized on-demand approach only provides services upon receiving clients’ requests. The existing proposal [2] is closest to our approach as our approach also falls into this category. However, it does not offer incentives for honesty, resistance to time source poisoning, or Sybil attack resistance. Incentives and Sybil attack resistance are resolved in the smart-contract-based proposals [8, 9, 10, 11, 12], but there lack of discussion on more secure time references of decryption time. More importantly, they also fall short in terms of reveal-verifiability under malicious clients, that is, malicious clients can construct some incorrect reveal-verification to frame an honest secret holder revealing incorrect secret share.
Presented in the remaining sections, our approach distinguishes itself by leveraging security features from smart contracts and incorporating secret sharing with reveal-verifiability under malicious clients, providing security features with flawless incentives that are not achieved in other existing approaches.
@@ II-B Secret Sharing Schemes for Secret Holders Approach
The original Shamir’s Secret Sharing scheme does not provide verifiability to the correctness of secret shares when sharing to the secret holders [43]; proposals on verifiable secret sharing [44, 45] provides verifiability in the distribution phase to ensure that holders receive correct shares.
However, it is not noticed in the existing work on the secret holder approach that verifiability in the distribution phase is not enough for timed-release cryptography. The message recovery stage where secret holders publish their secret shares is essential for timed-release cryptography systems, otherwise, it introduces attack opportunities for adversarial holders to submit incorrect shares to fail message recovery or for adversarial clients to make malicious claims that holders submit incorrect shares. To the best of our knowledge, secret sharing schemes with reveal-verifiability have not been demonstrated.
Moreover, secret sharing schemes often require peer-to-peer communication channels to distribute secret shares. When applied to timed-release cryptography, clients send secret shares to secret holders and also the time to reveal the shares. It is hard to ensure each secret holder receives the correct secret share and the same reveal time via a peer-to-peer channel. Applying existing secret sharing schemes in a public communication channel results in significant communication overhead, given that clients are required to publish all encrypted shares to deliver a unique secret share to each secret holder. This results in a large message size of $n\cdot|s|$, where $n$ is the number of holders and $|s|$ is the size of a secret share. Each holder receives all $n$ encrypted secret shares, even though only one share is relevant to each holder.
Based on the unique requirements of reveal-verifiability and public communication channels to apply secret sharing in timed-release cryptography, a new cryptographic protocol is introduced in this work to fulfill these two requirements.
@@ II-B1 Required Background
III Towards Secure Conditional Information Reveal Systems
Taking a step back from timed-release cryptography, this section introduces a paradigm to show that smart contract platforms can be applied to serve generic conditional information reveal systems with key security advantages.
III-A Security Challenges
While existing conditional information reveal systems exhibit certain similarities, their design is confronted by distinct challenges. For instance, as a widely used identity-based conditional information reveal system, a foundational presumption in prevailing single-sign-on (SSO) systems is the implicit trust required between service providers and clients towards authentication service providers. Consequently, its centralized architecture introduces vulnerabilities, most notably, a single point of failure. In the realm of location verification, it is challenging to identify a method that is simultaneously reliable, efficient, and precise for positioning. Despite the prevalence of conditional information reveal systems, a systematic approach to identifying the challenges they face is absent.
In summarizing the core components of conditional information reveal systems, we identify that the key security requirements for conditional information reveal systems include:
@@ III-A Security Challenges
III-B Decentralizing Conditional Information Reveal Systems
Considering the smart-contract-based proposals in time-based conditional information reveal system [8, 9, 10, 11, 12] in a generalized view, this architecture can in fact be applied to all types of conditional information reveal systems. The only difference is that the secret holders require different oracles to check if the required conditions are met.
In general, involving multiple secret holders through secret sharing and communicating through smart contracts improves system reliability. To elaborate, the advantages of such an architecture include:
Data authenticity and availability
Information holders and clients interact with smart contracts deployed on a blockchain network. This form of communication offers key advantages such as immutability and non-repudiation of messages, which are crucial for maintaining accountability and ensuring honest behavior among all parties involved in the system.
Fault tolerance and Sybil attack resistances
Involving multi-secret-holders with secret sharing protocols provides fault tolerance and robustness for the system, but it raises the risks of Sybil attacks in the permissionless setting where anyone can become a secret holder. Incorporating smart contracts for cryptocurrency staking addresses this issue.
Network autonomy through encoded rules in the smart contract
The smart contract provides a platform for secret holders or the wider public to raise disputes transparently. Network autonomy is achieved by disputing secret holders and using protocols for verifiable information sharing. For example, secret holders can dispute the actions of other holders via the smart contract. Sufficient disputes against a secret holder results in blacklisting and loss of cryptocurrencies as a punishment.
@@ III-B Decentralizing Conditional Information Reveal Systems
Figure 2 illustrates the modules of the smart-contract-based conditional information reveal architecture. Computation and verification of secret shares can be executed locally by holders to reduce on-chain computation costs; holders communicate and self-organize the network through a smart contract. This generic architecture provides adaptability to different conditional information reveal systems, accommodating various information sharing conditions.
IV A Smart-Contract-based Timed-Release Cryptography System
Going back into time-based conditional information reveal, this section illustrates a timed-release cryptography system using the decentralized conditional information reveal architecture. Sharing a similar basic idea with other smart-contract-based constructions [8, 9, 10, 11, 12], Our system additionally achieves higher reliability and security by making uses of the blockchain clock and achieving reveal-verifiability to prevent malicious clients from framing honest secret holders.
IV-A The Smart-Contract-based Architecture